# Gueter Josmy Faure1    Min-Hung Chen2    Jia-Fong Yeh1    Hung-Ting Su1 Shang-Hong Lai4    Winston H. Hsu1,3

### v3

@@ 4.1 Datasets and Evaluation Metrics
For long-form video classification, we utilize three diverse datasets. The first, LVU (Wu & Krahenbuhl, 2021), focuses on movie content, offering a rich source of narrative and thematic video data. The second, Breakfast (Tang et al., 2019), consists of instructional videos that emphasize procedural and step-by-step understanding. Lastly, COIN (Kuehne et al., 2014) is another instructional video dataset, but it covers a broader range of procedural activities compared to Breakfast. We report top-1 classification accuracy on these datasets.
For long-form video question answering, we employ the MovieChat-1k dataset (Song et al., 2024) and report both zero-shot and fully-supervised results. As evaluation metrics, we follow the evaluation protocol developed by Maaz et al. (2023), employing GPT-3.5-turbo (Brown et al., 2020) to assess both accuracy and answer quality score.
@@ 4.2 Quantitative Results
We present our long video action classification results in Table 1 for LVU (Wu & Krahenbuhl, 2021), Table 3 for Breakfast (Kuehne et al., 2014) and COIN (Tang et al., 2019), and compare HERMES ’s performance against transformer-based models, including Object Transformers (Wu & Krahenbuhl, 2021), Movies2Scenes (Chen et al., 2023), and FACT (Lu & Elhamifar, 2024); hybrid state-space and transformer-based models such as Vis4mer (Islam & Bertasius, 2022), TranS4mer (Islam et al., 2023), and S5 (Wang et al., 2023); as well as the LLM-based model MA-LMM (He et al., 2024). For the MovieChat-1k dataset (Song et al., 2024), our results are presented in Table 3, where we compare against recent LLM-based models including MovieChat (Song et al., 2024), Video-ChatGPT (Maaz et al., 2023), Video-LLaMA (Zhang et al., 2023), and VideoChat (Li et al., 2023b). Our method achieves state-of-the-art performance across all datasets, with notable accuracy improvements of 7.3% on LVU and 14.9% on MovieChat-1k, significantly surpassing previous methods.
Table 1: SOTA Comparison on the LVU Dataset: The table presents Top-1 accuracy for various models. Unlike the minor incremental improvements observed among other methods, our model demonstrates a significant performance leap, outperforming its nearest competitor by 7.3%. The highest score is highlighted in bold, and the second highest is underlined.
Model Content Metadata Avg
Relation Speak Scene Director Genre Writer Year
Object Transformer 53.1 39.4 56.9 52.1 54.6 34.5 39.1 47.1
VIS4mer 57.1 40.8 67.4 62.6 54.7 48.8 44.8 53.7
TranS4mer 59.5 39.2 70.9 63.9 55.9 46.9 45.5 54.5
S5 67.1 42.1 73.5 67.3 65.4 51.3 48.0 59.2
Movies2Scenes 71.2 42.2 68.2 70.9 57.8 55.9 53.7 60.0
MA-LMM 58.2 44.8 80.3 74.6 61.0 70.4 51.9 63.0
HERMES (Ours) 67.6 47.5 90.0 82.6 69.5 77.2 57.7 70.3
Table 2: Performance comparison on Breakfast and COIN datasets (Top-1 accuracy). Our method outperforms state-of-the-art models on both datasets.
Model Breakfast COIN
FACT 86.1 -
VIS4mer 88.2 88.4
MA-LMM 93.0 93.2
S5 90.7 90.8
TranS4mer 90.3 89.2
HERMES (Ours) 95.2 93.5
Model Global Breakpoint
Acc. Score Acc. Score
MovieChat 63.7 3.15 48.1 2.46
Video-ChatGPT 58.7 2.89 47.8 2.43
Video-LLaMA 56.3 2.72 45.8 2.11
VideoChat 60.2 3.08 46.3 2.32
HERMES (Ours) 78.6 4.23 57.3 3.29
HERMES (Ours) 84.9 4.40 65.8 3.65
4.3 Ablation Studies
Ablations are conducted on the MovieChat-1k test set (global mode) using the zero-shot setting with additional ablations on the Breakfast dataset using the fully-supervised setting. These experiments focus on our two primary contributions, ECO and SeTR. For an extended and more comprehensive ablations, please refer to Section A.4.
How important is ECO?
[…]
Table 4: Ablations on the memory update design of our Episodic COmpressor (ECO).
@@ 4.3 Ablation Studies
Table 5: Ablations on different semantic compression methods.
@@
Table 6: Performance comparison between frame Q-Former, video Q-Former and our hierarchical Q-Former architecture.
@@
Flat $fQFormer$ 93.2
Flat $vQFormer$ 94.1
Hierarchical $QFormer$ 95.2
Table 7: Zero-shot performance comparison of MA-LMM with and without ECO and SeTR  integration on MovieChat-1k.
Model Acc. Score Latency (s)
MA-LMM 73.3 4.05 467
MA-LMM + ECO 76.7 (+3.4) 4.14 (+0.09) 266 (-43%)
MA-LMM + SeTR 77.1 (+3.8) 4.16 (+0.11) 474 (+1.5%)
HERMES (Ours) 78.6 (+5.3) 4.23 (+0.18) 250 (-46%)
@@
How important is SeTR?
@@
Yes. We conducted an ablation study on the Breakfast dataset (Kuehne et al., 2014), to evaluate the efficacy of our proposed hierarchical Q-Former architecture. As shown in Table 4.3, our hierarchical Q-Former achieves superior performance with an accuracy of 95.2%, outperforming both flat frame-level ($fQFormer$, 93.2%) and video-level ($vQFormer$, 94.1%) architectures. This improvement can be attributed to the hierarchical structure’s ability to capture multi-scale features, effectively aggregating information from frame to video level. By first processing frame-level details and then aggregating them at the video level, our approach mitigates information loss that may occur in direct video-level processing while avoiding the computational intensity of processing every frame individually.
ECO as off-the-shelf memory manager.
We evaluate the impact of our Episodic COmpressor (ECO) as plug-ins to the existing MA-LMM model (He et al., 2024) by replacing the memory bank of MA-LMM. ECO is designed to efficiently process long video sequences while preserving temporal coherence and narrative structure. The results in Table 4.3 demonstrate substantial improvements with ECO integration with an accuracy increase by 3.4%. Moreover, ECO demonstrates superior efficiency compared to MA-LMM’s memory bank, almost halving the overall inference latency.
SeTR as off-the-shelf semantics retriever.
We also integrate our Semantics reTRiever (SeTR) into MA-LMM. SeTR is designed to enhance long video understanding by distilling high-level semantic cues, providing a cohesive framework for comprehending context and themes in long-form videos. As shown in Table 4.3, the integration of SeTR results in significant performance improvements. Accuracy increases by 3.8%, while the score improved by 0.11 points. Remarkably, these substantial performance gains were achieved with only a marginal 1.5% increase in inference time, highlighting the computational efficiency of SeTR and its potential for seamless integration into models requiring enhanced semantic representations.
Both SeTR and ECO yield significant performance enhancements when integrated with MA-LMM. SeTR showed a marginally higher performance boost in accuracy, which is expected given its role as a semantic add-on to MA-LMM’s existing memory management system.
How effective and efficient is ECO compared to other memory compressors?
@@
To demonstrate the effectiveness and efficiency of our proposed ECO, we conduct a comparative analysis against two strong existing compression techniques: ToMe (Bolya et al., 2022) and MA-LMM (He et al., 2024) in Figure 5. We calculate the inference time for each model on the MovieChat-1k dataset. Powered by ECO our model achieves the highest accuracy (78.6%) among all models, outperforming MA-LMM by 5.3% and ToMe by a substantial 13.8%. HERMES achieves the highest inference speed among the compared models, while also maintaining superior accuracy. It is slightly faster than ToMe and significantly outperforms MA-LMM, reducing inference time by 46% compared to the latter. These results demonstrate our model’s ability to deliver state-of-the-art accuracy without compromising on efficiency.
Hyperparameters for ECO and SeTR.
@@
Our experiments on the MovieChat-1k (zero-shot) and Breakfast (fully-supervised) datasets reveal compelling insights into the optimal configuration of ECO (Figure 4) and SeTR (Figure 4). For ECO, we discover that an episodic memory size of 20 consistently yields peak performance across both datasets, achieving a 78.6% accuracy on MovieChat-1k and a 95.2% on Breakfast. This sweet spot balances comprehensive video representation with computational efficiency, as larger memory sizes show diminishing returns. SeTR’s performance proved equally intriguing, with a keep ratio of 20% (reducing representations by 80%) emerging as the optimal choice. Remarkably, even at extreme keep ratios as low as 5%, our model exhibits robust performance, with a very slight decrease in accuracy. Such results demonstrate the resilience of HERMES to hyperparameter variations suggesting that it is suitable for deployment across diverse video understanding datasets with minimal hyperparameter tuning.
Figure 6: Qualitative Results: We select a challenging video from the MovieChat-1k dataset and pose various difficult questions to both MovieChat (Song et al., 2024) and HERMES. The results demonstrate our model’s superior ability to answer both fine-grained questions (Q1 and Q3) and general questions (Q2). Answers highlighted in blue denote tentative answers, red denote wrong answers, purple denote hallucinations, and green denote correct answers.
@@ 4.4 Qualitative Results
We present qualitative results on a challenging movie scene from the MovieChat-1k dataset to evaluate our model’s capability in answering both fine-grained and general questions about an extended video (14,000 frames). To rigorously assess the models, we bypass the original, basic Q&A from the dataset (e.g., Q: What’s the time in the video? A: Day; Q: Are there any shooting scenes? A: Yes) and pose questions that require a deeper understanding of the scene and actions, such as those concerning the expressions on the faces of the spectators. Our model demonstrates an ability to accurately respond to these questions while exhibiting a candid acknowledgment of its limitations (e.g., Q3). In contrast, MovieChat (Song et al., 2024) frequently generates hallucinated and incorrect answers. For instance, in response to Q2, MovieChat avoids answering the question directly and instead describes the scene. Notably, our model achieves this performance by processing only 100 out of the 14,000 frames (approximately 0.7%), whereas MovieChat processes 2,048 frames, which is more than 20 times the data utilized by our model. For additional qualitative results, including failure cases of our model, please refer to Section A.5.
5 Related Work
Action recognition is an essential task in video understanding, primarily focusing on identifying specific actions within short video clips. Various approaches have been developed, with convolutional neural networks (CNNs) forming the core of many of them. Early work by Ji et al. (2012) utilize 3D convolutions, while Varol et al. (2017) employed temporal convolutions. More recently, transformer-based models have gained prominence, as demonstrated in works such as Faure et al. (2023), Xu et al. (2021), and Zhang et al. (2022).
Video question answering (VideoQA) aims to answer questions related to video content, requiring a deep understanding of both visual and textual information. Datasets such as ActivityNet-QA (Yu et al., 2019) for short videos, and MovieChat-1k for long videos (Song et al., 2024) provide benchmarks for evaluating models in this field, allowing for several research endeavors on this subject (Zhang et al., 2020; Zhuang et al., 2020; Pan et al., 2023).
Long-form video understanding presents unique challenges due to the extended duration and complex narrative structures involved. Datasets with these properties include LVU (Wu & Krahenbuhl, 2021), COIN (Tang et al., 2019), Breakfast (Kuehne et al., 2014), and MovieChat-1k (Song et al., 2024). Traditional approaches to tackling such a task often extend methods designed for short videos to handle longer sequences namely pooling over the temporal dimension (Tang et al., 2020; Faure et al., 2023). Other methods such as Wu & Krahenbuhl (2021); He et al. (2024); Wu et al. (2022) and Song et al. (2024) explore memory techniques, emphasizing the need for more sophisticated models capable of managing large temporal spans. Wang et al. (2023) introduce selective structured state-spaces for long-form videos, followed by others Islam & Bertasius (2022); Islam et al. (2023) exploiting the ability of state-space models to retain long-term context.
LLM-based Long-Form Video Understanding: Recent advancements in large language models (LLMs) (Touvron et al., 2023; Chiang et al., 2023) have piqued researchers’ curiosity regarding their use for video understanding (Maaz et al., 2023). It turns out to be a good match, as understanding videos often involves transforming their content into words, whether it’s video captioning, video question answering, or even action classification. Song et al. (2024) and He et al. (2024) propose frameworks that employ memory techniques to handle extensive video content while Ren et al. (2024) presents TimeChat, explicitly conditioning the model to manage time-dependent information.
@@ 6 Conclusion
We propose HERMES, a novel framework designed to enhance long-form video understanding through two key components inspired by cognitive processes. The first, Episodic COmpressor (ECO), captures representations as sequences of continuous actions, reflecting episodic memory. The second, Semantics reTRiever (SeTR), serves as a high-level summarizer, distilling essential semantic information. Our model achieves state-of-the-art results on several long-video datasets, significantly outperforming existing methods. Through experiments on LVU, Breakfast, COIN and MovieChat, we have demonstrated the effectiveness and efficiency of ECO and SeTR.
Appendix A Appendix
A.1 Reproducibility Statement
To facilitate the reproducibility of our work, we will make our code, pretrained models, default hyperparameters, and preprocessed annotations publicly available. Detailed hyperparameters for each dataset are also provided in Table 8. Our model demonstrates efficient performance, completing inference on the MovieChat-1k test set in 13 minutes (22 FPS) using a single V100 GPU (32 GB), and training on the MovieChat-1k dataset in less than 12 minutes with 8x 32 GB GPUs. In contrast to recent LLM-based approaches that necessitate extensive and costly multi-stage pretraining on increasingly large datasets, our model is designed for accessibility, thereby lowering the barrier for researchers without access to high-end computing resources. We achieve high performance while maintaining accessibility by leveraging existing pretrained weights and implementing our training-free ECO and SeTR, resulting in a model where finetuning is optional.
A.2 Implementation Details
To ensure the reproducibility of our results, we provide the training details, which are also the defaults in our soon-to-be-released code. These settings are mostly consistent across different datasets.
Table 8: Hyperparameters used for different datasets.
@@ A.2 Implementation Details
LR is the learning rate, and Keep Ratio is the SeTR keep ratio. Episodes refer to the number of episodes to which we compress the input frames (i.e., the capacity of ECO). The number of frames (N) represents the quantity of frames retained from the original video to serve as input to the model. These frames are selected by applying a regular stride over the original video’s frame sequence, where the stride length is determined by the ratio of original frame count to N. Max Epoch = 20 means we run the program for 20 epochs, performing evaluation after each epoch, and then pick the model with the highest validation accuracy. MovieChat-1k (G) and MovieChat-1k (B) denote global and breakpoint modes, respectively. All models were trained on 8 V100 GPUs (32GB VRAM each).

### v4

@@ 5.1 Datasets and Evaluation Metrics
For long-form video classification, we utilize three datasets. The first, LVU [43], focuses on movie content, offering a rich source of narrative and thematic video data. The second, Breakfast [34], consists of instructional videos that emphasize procedural understanding. Lastly, COIN [18] is another instructional video dataset that covers a wider range of procedural activities compared to Breakfast. We report top-1 classification accuracy on these datasets.
For long-form video question answering, we employ the MovieChat-1k dataset [32] and report both zero-shot and fully-supervised results. As evaluation metrics, we follow the evaluation protocol developed by [22], employing GPT-3.5-turbo [4] to assess both accuracy and answer quality score. We also perform plug-and-play analysis of ECO and SeTR on three SOTA methods including MA-LMM [14], LongVA [51] and LLaVA-OneVision [19] and show enhanced performance on VideoMME [13] and MovieChat-1k.
@@ 5.2 Quantitative Results
For VQA, we evaluate on the MovieChat-1k dataset [32]. As shown in Table 1, HERMES surpasses recent LLM-based models including MovieChat [32], Video-ChatGPT [22], Video-LLaMA [49], and VideoChat [20], achieving a substantial 14.9% improvement over previous best results. On standard long-form video classification benchmarks including LVU [43], Breakfast [18], and COIN [34], HERMES consistently outperforms existing approaches (Table 2). We compare our model against three categories of methods: transformer-based models [43, 5, 21], hybrid architectures combining state-space and transformer approaches [15, 16, 41], and the LLM-based model MA-LMM [14]. Notably, HERMES achieves a 7.3% improvement over the previous state-of-the-art on LVU.
@@
Table 6: Ablations on the memory update design of our Episodic COmpressor.
@@
Table 7: Ablations on different semantic compression methods.
@@
$fQFormer$ 93.2
$vQFormer$ 94.1
$HQFormer$ 95.2
Table 8: Performance comparison between frame Q-Former, video Q-Former and our hierarchical Q-Former.
@@
5.3 Pilot Study: ECO and SeTR as plug-in modules
In this pilot study, we demonstrate the versatility and effectiveness of our Episodic COmpressor (ECO) and Semantics reTRiever (SeTR) by integrating them into three state-of-the-art video-language models: MA-LMM [14], LongVA [51], and LLaVA-OneVision [19] and evaluate them on two challenging benchmarks: MovieChat-1k [32] and VideoMME (w/o sub.) [13]. The results are reported in Tables 5, 3 and 4.
ECO: A Lightweight Episode Compressor. Integrating ECO, we notice consistent and substantial improvements across all models. Most notably, replacing MA-LMM’s memory bank with ECO yields a significant 3.4% increase in accuracy while simultaneously reducing inference time [^2] Inference time relative to the baseline by 43% and keeping memory usage constant. Similar efficiency gains are observed when integrated with LongVA [51] and LLaVA-OneVision [19], where ECO maintains or improves accuracy while reducing latency by 30% and 35%, respectively, and almost halving the GPU memory usage in the case of LongVA.
SeTR: An Efficient Semantics Retriever. To further validate the complementary nature of our modules, we integrate SeTRinto the same three models. As evidenced in Tables 5, 3, and 4, SeTR consistently enhances the models’ performance. For MA-LMM, we observe a substantial 3.8% increase in accuracy and a 0.11 improvement in score, achieved with only a minimal 1.5% increase in inference time. This demonstrates SeTR’s ability to extract rich semantic information while maintaining computational efficiency. When combined with ECO, the integration of SeTR into LongVA and LLaVA-OneVision yields further accuracy improvements of 0.37% for both models, while preserving the significant latency and memory usage reductions.
The consistent performance improvements achieved by both modules across different architectures and datasets underscore their effectiveness as plug-and-play solutions for enhancing video-language models. For MA-LMM, we evaluate SeTR in conjunction with their existing memory bank, demonstrating its ability to extract complementary semantic information. For LongVA and LLaVA-OneVision, we showcase the additive benefits of incorporating both modules sequentially, highlighting their synergistic relationship in improving model capabilities while maintaining efficiency.
5.4 Ablation Studies
Ablations are conducted on the MovieChat-1k test set (global mode) using the zero-shot setting with additional ablations on the Breakfast dataset using the fully-supervised setting. These experiments focus on our two primary contributions, ECO and SeTR. For extended and more comprehensive ablations, please refer to Section H.4 (in the Supp.). We also visualize the features extracted by each module in Section H.7.1
@@ 5.4 Ablation Studies
Do we need a hierachical Q-Former? Yes. We conducted an ablation study on the Breakfast dataset [18], to evaluate the efficacy of our proposed hierarchical Q-Former architecture. As shown in Table 8, our hierarchical Q-Former achieves superior performance with an accuracy of 95.2%, outperforming both flat frame-level ($fQFormer$) and video-level ($vQFormer$). This improvement can be attributed to the hierarchical structure’s ability to capture multi-scale features, effectively aggregating information from frame to video level. By first processing frame-level details and then aggregating them at the video level, our approach mitigates information loss that may occur in direct video-level processing while avoiding the computational intensity of processing every frame individually.
@@
How effective and efficient is ECO compared to other memory compressors? To demonstrate the effectiveness and efficiency of our proposed ECO, we conduct a comparative analysis against two strong existing compression techniques: ToMe [3] and MA-LMM [14] in Figure 5. We calculate the inference time for each model on the MovieChat-1k dataset. Powered by ECO, HERMES achieves the highest accuracy (78.6%) among all models, outperforming MA-LMM by 5.3% and ToMe by a substantial 13.8%. HERMES also achieves the highest inference speed among the compared models, while maintaining superior accuracy. It is slightly faster than ToMe and significantly outperforms MA-LMM, reducing inference time by 46% relative to the latter. These results demonstrate our model’s ability to deliver state-of-the-art accuracy without compromising on efficiency.
@@
Hyperparameters for ECO and SeTR. Our experiments on the MovieChat-1k (zero-shot) and Breakfast (fully-supervised) datasets reveal compelling insights into the optimal configuration of ECO (Figure 5) and SeTR (Figure 5). For ECO, we discover that an episodic memory size of 20 consistently yields peak performance across both datasets, achieving a 78.6% accuracy on MovieChat-1k and a 95.2% on Breakfast. This sweet spot balances comprehensive video representation with computational efficiency, as larger memory sizes show diminishing returns. SeTR’s performance proved equally intriguing, with a keep ratio of 20% (reducing representations by 80%) emerging as the optimal choice for both datasets. Such results demonstrate the resilience of HERMES to hyperparameter variations suggesting that it is suitable for deployment across diverse video understanding datasets with minimal hyperparameter tuning.
@@ 5.5 Qualitative Results
We present qualitative results on a challenging movie scene from the MovieChat-1k dataset (Figure 6) to evaluate our model’s capability in answering both fine-grained and general questions about an extended video (14k frames). To rigorously assess the models, we bypass the original Q&As from the dataset (e.g., Q: What’s the time in the video? A: Day, …) and ask questions that require a deeper understanding of the scene. Our model accurately responds to these questions while exhibiting a candid acknowledgment of its limitations (e.g., Q3). In contrast, MovieChat [32] frequently generates hallucinated and incorrect answers. HERMES achieves this performance by processing only 100 out of the 14k frames (approximately 0.7%), whereas MovieChat processes 2,048 frames, more than 20 times the data utilized by HERMES. We provide additional qualitative results and failure cases in the supplemetary material, Section H.7 and Section H.8.
Figure 6: Qualitative Results: We select a challenging video from the MovieChat-1k dataset and pose various difficult questions to both MovieChat [32] and HERMES. The results demonstrate our model’s superior ability to answer both fine-grained questions (Q1 and Q3) and general questions (Q2). Answers highlighted in blue denote tentative answers, red denote wrong answers, purple denote hallucinations, and green denote correct answers.
6 Limitations
While HERMES  demonstrates significant efficiency and performance gains, it relies on heuristics for both episodic compression and semantic retrieval, which may occasionally fail to capture subtle but important temporal details or contextual nuances. Furthermore, the episodic compressor and semantic retriever operate independently, potentially allowing redundancy. Due to computational constraints, we were unable to pretrain HERMES  on large-scale video datasets, limiting direct comparisons with extensively pretrained models like LLaVA-OneVision on benchmarks such as VideoMME. Nevertheless, the substantial improvements achieved through our lightweight integration approach suggest promising directions when combined with more computational resources.
@@ 7 Conclusion
We present HERMES, a method for enhancing long-form video understanding through two powerful, modular components inspired by cognitive processes. The Episodic COmpressor (ECO) captures representations as sequences of continuous actions while significantly reducing computational overhead, and the Semantics reTRiever (SeTR) serves as an efficient semantic enrichment mechanism. As standalone components, these modules can be seamlessly integrated into existing video-language models, consistently improving their performance while reducing inference latency. As a complete system, HERMES achieves state-of-the-art results across several long-video datasets, significantly outperforming existing methods. Through extensive experiments on five datasets, and integration studies with three SOTA models, we have demonstrated the effectiveness, efficiency, and versatility of our approach. As model sizes continue to increase and inference efficiency becomes a critical bottleneck in video understanding, our work provides a timely and foundational approach for both enhancing existing LLM-based systems and developing more scalable standalone solutions.
Supplementary Material
H Supplementary Material
H.1 Reproducibility Statement
To facilitate the reproducibility of our work, we will make our code, pretrained models, default hyperparameters, and preprocessed annotations publicly available. Detailed hyperparameters for each dataset are also provided in Table 9. Our model demonstrates efficient performance, completing inference on the MovieChat-1k test set in 13 minutes (22 FPS) using a single V100 GPU (32 GB), and training on the MovieChat-1k dataset in less than 12 minutes with 8x 32 GB GPUs. In contrast to recent LLM-based approaches that necessitate extensive and costly multi-stage pretraining on increasingly large datasets, our model is designed for accessibility, thereby lowering the barrier for researchers without access to high-end computing resources. We achieve high performance while maintaining accessibility by leveraging existing pretrained weights and implementing our training-free ECO and SeTR, resulting in a model where finetuning is optional. We also demonstrate the applicability of our modules to existing video models, and are planning to submit pull requests to integrate our modules into these models.
For fully-supervised results, QFormers and adapter are fine-tuned on the respective dataset’s training split. For plug-in experiments, ECO and SeTR are inserted into target architectures at inference time, with zero additional training, demonstrating true plug-and-play capability.
@@ H.1 Reproducibility Statement
VideoMME (LongVA) - - 1 128 32 0.125
VideoMME (Llava-OV) - - 1 128 32 0.125
Table 9: Hyperparameters used for different datasets.
H.2 Implementation Details
To ensure the reproducibility of our results, we provide training and inference details in Table 9. These settings are mostly consistent across different datasets. In the table, LR is the learning rate, and Keep Ratio is the SeTR keep ratio. Episodes refer to the number of episodes to which we compress the input frames (i.e., the capacity of ECO). The number of frames (N) represents the quantity of frames retained from the original video to serve as input to the model. These frames are selected by applying a regular stride over the original video’s frame sequence, where the stride length is determined by the ratio of original frame count to N. Max Epoch = 20 means we run the program for 20 epochs, performing evaluation after each epoch, and then pick the model with the highest validation accuracy. MovieChat-1k (G) and MovieChat-1k (B) denote global and breakpoint modes, respectively. All models were trained on 8 V100 GPUs (32GB VRAM each). We test on VideoMME using the zero-shot setting by applying our modules to two different models, the same parameters were used across models for consistency.